# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 2.2.2. Workflow Authoring
The second dimension is the presence or absence of a Predefined Workflow, that is a set of steps conducted to repair issues.
(1) Workflow authored by humans: The overall flow is predefined by humans. Traditional APR systems such as GenProg  (legoues2012genprogtse, ) usually implement a two or three- main steps workflows, which are executed sequentially: First a Fault Localization stage followed by Patch Generation stage, and finally Patch validation stage. That workflow usually is hard-coded in the approach implementation. Note that inside one of the workflow steps, autonomy may be present, nevertheless the control structure of the principal workflow is fixed. (2) Emergent workflow: There is not predefined workflow crafted by humans. It corresponds to fully agentic systems which are driven entirely by one or more autonomous agents. The execution path is dynamic and reactive, based on goals, environment or feedback.
@@ 2.2.3. Autonomy over execution path: Control flow autonomy
    Scaffolded Execution: A human provides a structured workflow, including stages (e.g., issue localization, patch generation, validation) and allowed transitions between them. The system is free to make decisions within and across these stages. This structure allows local autonomy within a globally defined flow.
@@
2.2.4. LLM-based Agents for End-to-end Software Maintenance
We follow the taxonomy of LLM-based agents for software engineering task by Liu et al. (liu2024SurveyAgentsLLM4SE, ). In particular, we focus on the pipeline of LLM-based Agents for End-to-end Software Maintenance that the authors present, which is composed of the following phases: a) Preprocessing, b) Issue Reproduction, c) Issue Localization, d) Task Decomposition, e) Patch Generation, f) Patch Verification, g) Ranking. For each of these phases, we discuss how the approaches face them. Even those not all the solutions are based on agents, the phases are also present in “traditional” APR approaches such as the Generate-and-Validate (LeGoues2011genprog, ), which conduct three of these phases: a) Bug localization (related to the mentioned Issue Localization), b) Patch Generation, c) Patch Verification.
@@ 2.3. Collected Data
DescriptionOrigin Scientific Articles Blog Github Linkedin Nothing SWECommit Whitepaper
Leaderboard Lite Verified Lite Verified Lite Lite Verified Lite Verified Lite Verified Lite
Total 31 27 12 24 4 1 2 3 2 10 13 1
Table 1. Count of Artifacts Describing the Approach and/or Experiment Conducted on SWE-Bench.
Table 1 presents the distribution of artifact types used to describe each leaderboard entry. In this paper, we analyze these artifacts, together with the information available in the SWE-Bench leaderboars -available online, to answer the research questions. Let us discuss the distribution of the type artifacts that we analyze. In both the SWE-Bench Lite and SWE-Bench Verified leaderboards, the most common artifact type is scientific articles (including preprints on arXiv), accounting for 31 and 27 entries, respectively. The scientific venues and journals where these articles have been published include both Software Engineering and AI/Machine Learning conferences. In the Software Engineering field, we identified publications at ICSE 2025 (ruan2024specrovercodeintentextraction, ), FSE 2025 (xia2024agentlessdemystifyingllmbasedsoftware, ; ma2025alibabalingmaagent, ) and ISSTA 2024 (Zhang2024AutoCoderRover, ). In the AI and Machine Learning domain, relevant articles were published at ICLR 2024 (wang2025solvedSWEBench, ), ICLR 2025 (ouyang2024repographenhancingaisoftware, ; wang2024openhandsopenplatformai, ), ICML 2025 (zainullina2025guidedsearchstrategies, ) and NeurIPS 2024 (yang2024sweagentagent, ).
Blog posts represent the second most frequent type of artifact used to describe the experiments conducted on SWE-Bench, to announce results, and to explain the proposed solutions. We associated 12 and 24 entries to blog posts in SWE-Bench Lite and SWE-Bench Verified, respectively. There is also one entry, CodeFixer from Globant, where the description was provided in a white paper available on the company’s website, and was not published on arXiv or any other pre-print repository at the time of writing this article. Moreover, some entries provide descriptions of the experiments and/or approaches in the README.me files included in the commits to SWE-Bench that introduce the experimental results. This is the case for 10 and 13 submissions in SWE-Bench Lite and SWE-Bench Verified, respectively. For example, the submission named SIMA mentions in the README.md file  [^10] https://github.com/SWE-bench/experiments/tree/main/evaluation/lite/20240706_sima_gpt4o that the implementation is based on the Agent Forest method presented by Li et al. (li2024MoreAgentsneed, ).
@@ 3. Results
3.1. RQ : What are the key characteristics of the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards across dimensions such as submitter origin, product type, availability, and LLMs used?
3.1.1. Analysis of Solutions
Descriptive statistics.
By May 25th 2025, we count, in total, 68 and 79 entries on SWE-Bench Lite and SWE-Bench Verified, respectively. As Fig. 1 and Table 2 show, this small margin in favor of SWE-Bench Verified comes mainly from a larger number of industry submissions, increasing from 44% to 65%; especially notable is the case of large publicly traded companies, with 32% over the total of solutions in SWE-Bench Verified compared to 10% in SWE-Bench Lite. The absolute number of academic submissions remains the same, and open source community solutions virtually disappear in SWE-Bench Verified, the same as single developer solutions. Solutions from small companies prevail in both leaderboards, and together with solutions from academia, they represent half of submissions not only in total numbers, but also individually in each leaderboard.
There are 26 out of the 147 entries (18%) submitted to both leaderboards. More than half (15) come from academia. If we count these repeated submissions only once, the total percentage of academy solutions decreases from 25% to 17% or, conversely, the percentage of solutions submitted to only one leaderboard by academia does not exceed 7%.
@@ 3.1.1. Analysis of Solutions
Academia (AC) 16 3 18.16 40.67 16 3 22.8 62.4 32 3
@@
Collaboration Ac-Ind (CAI) 9 8 30 47 4 4 56.6 64.6 13 10
Collaboration Academia (CA) 2 1 24 24.67 2 1 31.5 32.8 4 1
@@
Company-Large (CL) 4 2 34.84 39.33 2 1 58.2 66.4 6 2
Company-Large-Publicly Traded (CLP) 7 5 27.33 48.33 18 7 39.2 68.2 25 9
Company-Medium (CM) 1 1 49 49 5 3 62.8 64.6 6 3
Company-Small (CS) 17 13 37 55 25 14 48.6 65.8 42 20
Company-Unknown (CU) 1 1 44.67 44.67 - - - - 1 1
OpenSource Community (OSC) 6 2 28.68 39 1 1 63.4 63.4 7 3
@@
SWE-Bench Lite SWE-Bench Verified Total
Submitter Type % Resolved % Resolved
#E #S Median Max #E #S Median Max #E #S
Academia (AC+CA) 18 4 20.7 40.7 18 4 26.7 62.4 36 4
Industry (CL+CLP+CM+CS+CU+CI) 30 22 35.7 55.0 51 26 48.6 68.2 81 36
Academia-Industry (CAI+AS) 11 9 30.0 47.0 8 6 48.9 64.6 19 12
Open-Source (OSC) 6 2 28.7 39.0 1 1 63.4 63.4 7 3
Single Developer (SD) 2 2 29.0 39.3 - - - - 2 2
Unknown (UNK) 1 1 42.7 42.7 1 1 41.6 41.6 2 2
Table 3. Descriptive statistics of submitter organization categories. #E shows the total number of entries, #S is the number of distinct submitters.
Figure 1. Solutions by category of submitter: accumulated (left), SWE-Bench Lite (center) and SWE-Bench Verified (right).
Temporal analysis. Fig. 2(a) shows the cumulative number of entries in each leaderboard, taking the dates from the leaderboards themselves. The vertical dotted red line shows the release date of SWE-Bench Verified (August 13th 2024), considered to be the date in which OpenAI released the blog article that describes the leaderboard [^11] https://openai.com/index/introducing-swe-bench-verified/. We observe that there are entries whose date precedes that official release. The reason of this is that the mentioned article from OpenAI also included, in addition to well defined construction criteria of SWE-Bench Verified, the results of an evaluation of open-source approaches (open-source scaffolds according to that article) such as AutoCoderRover, Agentless and SWE-Agent on SWE-Bench Verified, which were already present in SWE-Bench Lite. The SWE-Bench Verified leaderboard includes these entries with the same date as SWE-Bench Lite. [^12] Added in the leaderboard by this commit: https://github.com/SWE-bench/swe-bench.github.io/commit/8d66692d5f55101c3c0f53daac3f8a479d29406e For example, the entry SWE-agent + GPT 4 (1106) is present in both leaderboards with the date 2024-04-02 (a date before the release of SWE-Bench Verified).
The figure shows how the leaderboards have evolved along five stages:
  1. (1)
    Nonexistence of SWE-Bench Verified. From Oct. 2023 until May 2024, all models were in SWE-Bench Lite only.
  2. (2)
    Prevalence of SWE-Bench Lite. Until September 2024, both repositories co-existed but still the growth rate of SWE-Bench Lite was greater than that of SWE-Bench Verified’s.
  3. (3)
    Prevalence of SWE-Bench Verified. In only two months, until Nov. 2024, SWE-Bench Verified experienced a sudden growth until reaching the same size as SWE-Bench Lite (41 entries each).
  4. (4)
    Similarity of both leaderboards. Until Feb. 2025, the number of entries remained approximately the same for both SWE-Bench Lite and SWE-Bench Verified.
  5. (5)
    Abandonment of SWE-Bench Lite. While submissions to SWE-Bench Verified continue to grow, those to SWE-Bench Lite appear to have plateaued.
(a) Entries by date. The vertical line correspond to the official release of SWE-Bench Verified by OpenAI.
@@ 3.1.2. Analysis of Submitter Organization
SWE-Bench Lite SWE-Bench Verified
% Resolved % Resolved
Submitter ST #E Median Max Submitter ST #E Median Max
Princeton University AC 11 4.3 28 Princeton University AC 12 12.6 62.4
Moatless OSC 5 30.7 39 EPAM CLP 5 39.6 62.8
Univ.IllinoisUC AC 4 32.2 40.7 Alibaba CLP 4 21.6 28.8
Bytedance CL 3 34 39.3 nFactorial CS 4 36.6 49.2
Amazon CLP 2 25 29.7 Amazon CLP 4 46.9 65.4
AppMap CS 2 28.8 36 Anthropic CM 3 49 63.2
All Hands et al. CAI 2 34.2 41.7 Univ.IllinoisUC AC 3 42.4 50.8
Isoform CS 2 45 55 Solver AI CS 3 45.4 50
Shanghai AI et al. AC 2 24 24.7 AutoCodeRover AS 3 46.2 51.6
OpenCSG CS 2 31.7 39.7 All-Hands CS 2 63.3 65.8
Table 4. Top-10 Submitters on each leaderboard. ST is submitter type, #E shows the number of total entries.
Figure 3. Organizations per leaderboard. In parenthesis the number of submissions: for those organization with submissions to both leaderboard we append two numbers $(l,v)$: $l$ is the submissions to Lite and $v$ to Verified.
@@
Tables 2 and 3and Fig. 4 show the results. There are 59 submitters in total, with more submitters to SWE-Bench Lite than to SWE-Bench Verified (40 submitters vs. 38). Therefore, the ratio of submissions per submitter to SWE-Bench Verified is greater than that of SWE-Bench Lite (2.07 vs. 1.70). Most submitters are companies, up to 61% in total, growing up to 81% if we also count industry-academy collaborations.
Figure 4. Submitters per category: total (left), SWE-Bench Lite (center) and SWE-Bench Verified (right).
The case of SWE-Bench Lite. The dominance of companies is not as high as in SWE-Bench Verified but still more than half (55%), and from them, small companies prevail (up to 13 submitters, representing 32.5% over the total). Out of these 13 small companies, 12 are still active (i.e., releasing products or posting updates in the social networks such as LinkedIn). We also found two entries submitted by an Academia spin-off, AutoCodeRover, which has been acquired by an established company, Sonar, in Feb. 2025.131313https://www.sonarsource.com/company/press-releases/sonar-acquires-autocoderover-to-supercharge-developers-with-ai-agents/ (Last access Feb. 25, 2025)
Large companies, including publicly traded ones, and medium companies too, have also submitted their solutions. It is worth noting that only one of them, Bytedance, has submitted more than two solutions to SWE-Bench Lite. Other big players such as IBM, Microsoft, Amazon and Alibaba have submitted only one or two. Some others have not submitted any separately, but they have done through some industry-academia collaboration: Google (su2025learnbyinteract, ), Intel (yu2025orcaLocallmAgent, ) and Meta (wei2025SWE-RL, ). Other collaborations have involved small companies founded in recent years such as All-Hands (wang2024openhandsopenplatformai, ).
@@
Four submitters were from academia, one of them being Princeton University, the founder of SWE-Bench. Another university from USA (University of Illinois Urbana-Champaign) and two Chinese academic submitters (Peking University and a collaboration among four research centers) complete the academic landscape. We also found entries submitted by two open source communities (Moatless and Aider). Last, there are two submissions done by single developers: (i) SIMA, an experiment aimed at testing the multi-agent architectures proposed by Li et al. (li2024MoreAgentsneed, ); (ii) Aegis, a “side project that I’m using as a learning opportunity to get hands-on with applications of agents for software engineering”. It is important to note that these two entries from single developers were built on top of the open-source Moatless framework, highlighting the importance of providing reusable artifacts and evaluation infrastructure, as Moatless does.
@@
The case of SWE-Bench Verified. As commented, the prevalence of industry in SWE-Bench Verified increases significantly, up to 68.4%. Per type of companies, fluctuations are minor; we may remark though the involvement of NVIDIA, Google and EPAM. Concerning academia, the same four submitters to SWE-Bench Lite contribute to SWE-Bench Verified; however, the number of industry-academy collaborations drop by the half (from 8 to 4).

### v2

@@ 2.2.1. Workflow Authoring
The first dimension is the presence or absence of a predefined workflow, that is, a set of steps conducted to repair issues.
(1) Workflow authored by humans: The overall flow is predefined by humans. Traditional APR systems such as GenProg  (legoues2012genprogtse, ) usually implement a workflow consisting of two or three main steps, which are executed sequentially: first, a Fault Localization stage, followed by Patch Generation stage, and finally Patch validation stage. That workflow is usually hard-coded in the approach implementation. Note that inside one of the workflow steps, autonomy (see next dimension) may be present, nevertheless the control structure of the principal workflow is fixed. (2) Emergent workflow: There is no predefined workflow crafted by humans. It corresponds to fully agentic systems which are driven entirely by one or more autonomous agents. The execution path is dynamic and reactive, based on goals, environment or feedback.
@@ 2.2.2. Autonomy over execution path: Control flow autonomy
    Scaffolded Execution: A human provides a structured workflow, including stages (e.g., issue localization, patch generation, validation) and allowed transitions between them. The system —typically leveraging an LLM— is free to make decisions within and across these stages. This structure -known as a Scaffold- allows local autonomy within a globally defined flow.
@@
2.2.3. Number of Agents
A quick observation of the leaderboard entries reveals, at least by name, that most of the solutions are based on Agents (e.g., SWE-Agent). Nevertheless, we define the following categories: a) No Agent. This includes repair systems that just prompt LLMs, e.g., using Chain of Thoughts; b) Single Agent; c) Multiple Agents.
Please note that some combinations are not allowed, e.g. an approach cannot have an emergent workflow and be non-agentic.
2.3. LLM-based Agents for End-to-end Software Maintenance
For responding RQ3, we follow the taxonomy of LLM-based agents for software engineering task by Liu et al. (liu2024SurveyAgentsLLM4SE, ). In particular, we focus on the pipeline of LLM-based Agents for End-to-end Software Maintenance that the authors present, which is composed of the following phases: a) Preprocessing, b) Issue Reproduction, c) Issue Localization, d) Task Decomposition, e) Patch Generation, f) Patch Verification, g) Ranking. For each of these phases, we discuss how the approaches face them. Even those not all the solutions are based on agents, the phases are also present in “traditional” APR approaches such as the Generate-and-Validate (LeGoues2011genprog, ), which conduct three of these phases: a) Bug localization (related to the mentioned Issue Localization), b) Patch Generation, c) Patch Verification.
@@ 2.4. Collected Data
Type Scientific Articles Blog SWECommit Github Linkedin Whitepaper Nothing
Leaderboard Lite Verified Lite Verified Lite Verified Lite Verified Lite Verified Lite Verified Lite Verified
Total 39 40 18 38 10 6 6 0 1 1 1 0 3 3
Table 1. Distribution of approach documentation types across the SWE-Bench Lite and SWE-Bench Verified leaderboards. Each column shows the number of entries associated with a specific documentation source.
We collected the entries published in the leaderboards until July 17th. [^10] Version of SWE-Bench studied: https://github.com/SWE-bench/swe-bench.github.io/commit/c53e6f9b8ea0b17060ef3a7718d3020e529f8a22 Table 1 presents the distribution of artifact types used to describe each leaderboard entry. In this study, we analyze these artifacts, together with the information available in the SWE-Bench leaderboards -available online, to answer the research questions. Let us discuss the distribution of the type artifacts that we analyze. In both the SWE-Bench Lite and SWE-Bench Verified leaderboards, the most common artifact type is scientific articles (including preprints on arXiv), accounting for 39 and 40 entries, respectively. The scientific venues and journals where these articles have been published include both Software Engineering and AI/Machine Learning conferences. In the Software Engineering field, we identified publications at ICSE 2025 (ruan2024specrovercodeintentextraction, ), FSE 2025 (xia2024agentlessdemystifyingllmbasedsoftware, ; ma2025alibabalingmaagent, ) and ISSTA 2024 (Zhang2024AutoCoderRover, ) and 2025 (ma2024lingmaswegpt, ). In the AI and Machine Learning domain, relevant articles were published at ICLR 2024 (wang2025solvedSWEBench, ), ICLR 2025 (ouyang2024repographenhancingaisoftware, ; wang2024openhandsopenplatformai, ), ICML 2025 (zainullina2025guidedsearchstrategies, ; li2025patchpilot, ; sohrabizadeh2025nemotroncortexa, ) and NeurIPS 2024 (yang2024sweagentagent, ).
Blog posts represent the second most frequent type of artifact used to describe the experiments conducted on SWE-Bench, to announce results, and to explain the proposed solutions. We associated 18 and 38 entries to blog posts in SWE-Bench Lite and SWE-Bench Verified, respectively. There is also one entry, CodeFixer from Globant, where the description was provided in a white paper available on the company’s website, and was not published on arXiv or any other pre-print repository at the time of writing this article. Moreover, some entries provide descriptions of the experiments and/or approaches in the README.me files included in the commits to SWE-Bench that introduce the experimental results. This is the case for 10 and 6 submissions in SWE-Bench Lite and SWE-Bench Verified, respectively. We also detect entries (3 on each leaderboard) not related to any artifact describing the experiment or the approach.
@@ 3. Results
3.1. RQ1: What are the key characteristics of the submissions to the SWE-Bench Lite and SWE-Bench Verified leaderboards across dimensions such as submitter origin, product type, availability, and LLMs used?
We begin with an overview of the number of entries and distinct approaches submitted to the leaderboards. Then, we analyze the approaches along four dimensions: (1) submitter profile, (2) product type, (3) availability, and (4) underlying LLM(s) used.
3.1.1. Overview and Temporal Analysis of SWE-Bench Leaderboard Submissions
By July 17th 2025, we count, in total, 79 and 99 entries on SWE-Bench Lite and SWE-Bench Verified, respectively. We identified 52 and 50 distinct approaches, respectively, including Agentless (xia2024agentlessdemystifyingllmbasedsoftware, ) and AutoCodeRover (Zhang2024AutoCoderRover, ). Among these, 30 approaches appear only in SWE-Bench Lite, 28 only in SWE-Bench Verified, and 22 in both. The complete list of approaches and submissions is provided in the appendix (Apprendix, ).
Fig. 1(a) shows the cumulative number of entries on each leaderboard, taking the dates from the leaderboards themselves. The vertical dotted red line shows the release date of SWE-Bench Verified (August 13th 2024), considered to be the date in which OpenAI released the blog article that describes the leaderboard [^11] https://openai.com/index/introducing-swe-bench-verified/. We observe that there are entries whose date precedes that official release, being the reason that the mentioned article also included, in addition to well defined construction criteria of SWE-Bench Verified, the results of an evaluation of open-source approaches (open-source scaffolds according to that article) such as AutoCoderRover, Agentless and SWE-Agent on SWE-Bench Verified, which were already present in SWE-Bench Lite. The SWE-Bench Verified leaderboard includes these entries with the same date as SWE-Bench Lite. [^12] Added in the leaderboard by this commit: https://github.com/SWE-bench/swe-bench.github.io/commit/8d66692d5f55101c3c0f53daac3f8a479d29406e For example, the entry SWE-agent + GPT 4 (1106) is present in both leaderboards with the date 2024-04-02 (a date earlier than SWE-Bench Verified’s release).
The figure shows how the leaderboards have evolved along five stages:
  1. (1)
    Nonexistence of SWE-Bench Verified. From Oct. 2023 until May 2024, all submission were in SWE-Bench Lite only.
  2. (2)
    Prevalence of SWE-Bench Lite. Until mid-September 2024, both repositories co-existed but still the growth rate of SWE-Bench Lite was greater than that of SWE-Bench Verified’s.
  3. (3)
    Prevalence of SWE-Bench Verified. In only two months, until mid-Nov. 2024, SWE-Bench Verified experienced a sudden growth until reaching the same size as SWE-Bench Lite (41 entries each).
  4. (4)
    Similarity of both leaderboards. Until mid-Feb. 2025, the number of entries remained approximately the same for both SWE-Bench Lite and SWE-Bench Verified.
  5. (5)
    Reduced activity on SWE-Bench Lite. While submissions to SWE-Bench Verified continue to grow steadily, activity on SWE-Bench Lite has significantly slowed, with only a few recent entries—such as those involving Claude 4 LLM— being added.
Figure 1(b) shows the evolution of the % Resolved metric, which is the primary measure reported in both leaderboards. Despite a period among Oct. 2024 and Nov. 2024 in which some solutions uploaded to SWE-Bench Verified were not optimal, the precision is notably higher in SWE-Bench Verified than in SWE-Bench Lite. Specifically, the median and maximum precision are 46.9% and 75.2% for SWE-Bench Verified, compared to 31.5% and 60.0% for SWE-Bench Lite. This difference may be attributed to the construction of SWE-Bench Verified, which was derived from the original SWE-Bench by manually filtering out instances considered too hard or unsolvable, with the aim of increasing the benchmark’s robustness and reliability (chowdhury2024swebenchverified, ).
(a) Entries by date. The vertical dotted line correspond to the official release of SWE-Bench Verified by OpenAI.
Figure 1.
@@ 3.1.1. Overview and Temporal Analysis of SWE-Bench Leaderboard Submissions
Single Academy (SA) 19 4 23 56.67 17 3 23.2 66.6 36 4
Collaboration Academia (CA) 6 3 41.34 60.33 3 2 32.8 46 9 4
Total Academia 25 7 27.33 60.33 20 5 31.5 66.6 45 8
@@
Collaboration Ac-Ind (CAI) 10 9 29.84 47 7 6 53 64.6 17 12
Total Academia-Industry 12 10 29.835 47 11 8 46.2 64.6 23 14
Company-Large (CL) 4 2 34.84 39.33 4 1 68.5 75.2 8 2
Company-Large-Publicly Traded (CLP) 7 5 27.33 48.33 21 8 39.6 68.2 28 10
Company-Medium (CM) 1 1 49 49 8 4 63.9 73.2 9 4
Company-Small (CS) 19 15 38 60 31 16 51.8 74.4 50 23
Company-Unknown (CU) 1 1 44.67 44.67 - - - - 1 1
@@
Total Industry 32 24 35.8 60 65 30 50 75.2 97 41
OpenSource Community (OSC) 6 2 28.68 39 2 2 67.1 70.8 8 3
@@
(a) Accumulated
(b) SWE-Bench Lite
(c) SWE-Bench Verified
Figure 2. Total Entries by category of submitter: accumulated (left), (center) and SWE-Bench Verified (right).
@@ 3.1.2. Analysis of Submitter Organization
Figure 3. Organizations per leaderboard. In parenthesis the number of submissions: for those organization with submissions to both leaderboard we append two numbers $(l,v)$: $l$ is the submissions to SWE-Bench Lite and $v$ to SWE-Bench Verified.
@@
Figure 3 presents the organizations that submitted results to the leaderboards (referred to as submitters), grouped by submitter type (e.g., Industry, Academia). The figure displays the name of each submitter along with the number of submissions (i.e., entries) made to each leaderboard. Some submitter names have been abbreviated, particularly in the case of collaborations involving multiple organizations. The full names are provided in the Appendix (Apprendix, ). There are 71 distinct submitters in total, with similar number of submitters to SWE-Bench Lite and to SWE-Bench Verified (47 submitters vs. 46). Therefore, the ratio of submissions per submitter to SWE-Bench Verified is greater than that of SWE-Bench Lite (2.15 vs. 1.68).
Table 2 summarizes, for each submitter type, the number of entries (#E), the number of distinct submitters (#S), and the median and maximum % Resolved scores. The categories Industry and Academia are further broken down into sub-types, as described in Section 2.1.2. As also illustrated in Figure 4, the majority of submitters are companies, accounting for 58% of the total. This proportion rises to 78% when including industry–academia collaborations. We now examine the characteristics of each leaderboard individually.
The difference in the number of entries between SWE-Bench Lite (79) and SWE-Bench Verified (99) comes mainly from a larger number of industry submissions, as Figure 2 shows, increasing from 41% (32 entries) to 66% (65 entries). Especially notable is the increase in submissions from large publicly traded companies, which account for 21.2% (21 entries) of all entries in SWE-Bench Verified, compared to 8.8% (7 entries) in SWE-Bench Lite. The absolute number of academic submissions remains on a similar level (25 vs 20), and open source community solutions virtually disappear in SWE-Bench Verified, the same as single developer solutions. Solutions from small companies prevail in both leaderboards, and together with solutions from academia, they represent half of submissions not only in total numbers, but also individually in each leaderboard.
(a) Accumulated
(b) SWE-Bench Lite
(c) SWE-Bench Verified
Figure 4. Total Distinct Submitters by category of submitter: accumulated (left), (center) and SWE-Bench Verified (right).
The case of SWE-Bench Lite. The dominance of companies between the submitters is not as high as in SWE-Bench Verified but still more than half (51%), and from them, small companies prevail (up to 15 submitters, representing 32.6% over the total). We also found two entries submitted by an Academia spin-off, AutoCodeRover, which has been acquired by an established company, Sonar, in Feb. 2025.131313https://www.sonarsource.com/company/press-releases/sonar-acquires-autocoderover-to-supercharge-developers-with-ai-agents/ (Last access Feb. 25, 2025)
Large companies—including publicly traded and medium-sized firms—have also submitted their solutions, typically independently and without external collaboration (e.g., IBM, Amazon). In contrast, some companies submitted exclusively through collaborative efforts, primarily with academic partners: Google (su2025learnbyinteract, ), Intel (yu2025orcaLocallmAgent, ) and Meta (wei2025SWE-RL, ). Other collaborations have involved small companies founded in recent years such as All-Hands (wang2024openhandsopenplatformai, ).
@@
Seven submitters to SWE-Bench Lite are from academia, one of them being Princeton University, the founder of SWE-Bench. Another two universities from the USA (University of Illinois Urbana-Champaign and Columbia University) and four Chinese academic submitters (Peking University and collaboration among Chinese universities and research centers) complete the academic landscape. We also found entries submitted by two open source communities (Moatless and Aider). Last, there are two submissions done by single developers: (i) SIMA, an experiment aimed at testing the multi-agent architectures proposed by Li et al. (li2024MoreAgentsneed, ); (ii) Aegis, a “side project that I’m using as a learning opportunity to get hands-on with applications of agents for software engineering”. It is important to note that these two entries from single developers were built on top of the open-source Moatless framework, highlighting the importance of providing reusable artifacts and evaluation infrastructure, as Moatless does.
@@
The case of SWE-Bench Verified. As commented, the prevalence of industry in SWE-Bench Verified increases significantly: the 65% of the submitters and the 66% of the entries are from industry. Per type of companies, fluctuations are minor; we may remark, though, the involvement of NVIDIA, Google, Anthropic and EPAM. Concerning academia, a new collaboration between two U.S. universities (UCSB and Columbia) contributed to SWE-Bench Verified.